OpenStudy (sugarycola):

Help please!!! I will give medal and become a fan :) questions will be in comment section help is appreciated

OpenStudy (sugarycola):

Which equation has intercepts at X(1, 0, 0), Y(0, 1, 0), and Z(0, 0, 2)? x y + 2z = 4 3x 2y + 6z = 12 2x 2y + z = 2 4x 4y + z = 4

OpenStudy (sugarycola):

Which equation is equivalent to x + 3y z = 3? 3x + 6y 3z = 6 4x + 12y 4z = 12 2x + 4y 2z = 8 2x + 3y 4z = 12

OpenStudy (sugarycola):

Which of the following points lies in the plane 3x 2y + 4z = 12? (2,2, 3) (5, 15, 2) (4, 16, 2) (4, 3,2)

OpenStudy (anonymous):

1.c

OpenStudy (anonymous):

u got answer?